# MPTCP in Single Radio Access Networks: a Paradox or an Opportunity?

### v1

@@ Abstract
This paper addresses the use of MultiPath Transmission Control Protocol (MPTCP) in a single Radio Access Technology (RAT) network. Different from other studies where multiple network access technologies are explored by the MPTCP, in this work we assess and evaluate the capability of MPTCP to operate over a single radio access network environment. With a vehicular network as use case, we show how the IEEE 802.11p interface is shared among the multiple logical links created between the On-Board Unit (OBU) and the several Road Side Units (RSUs) in its range, supporting the different MPTCP subflows. The results, obtained through experimentation with real vehicular networking hardware, show that MPTCP allows for seamless handovers, ensuring continuous, stable and efficient communication in highly mobile environments.
@@
Vehicular Networks, MultiPath Transmission Control Protocol, Single Radio Access Networks, IEEE 802.11p.
@@ I Introduction
MPTCP is an enhanced version of Transmission Control Protocol (TCP) designed to increase application performance and reliability by allowing concurrent use of multiple paths between the source and the destination [1]. A fundamental aspect of MPTCP is its resilience in scenarios of heterogeneous access technologies, exploiting the diversity of access technologies, dynamically distributing data across multiple paths. This ensures efficient and robust communication, allowing the protocol to adapt to the various conditions and capabilities of each network. This protocol also ensures seamless handovers between various network interfaces, such as Wi-Fi and cellular networks, a feature particularly beneficial in mobile environments, characterized by frequent changes in the network topology.
@@
This work presents a different study regarding the exploration and performance of MPTCP in VANETs. Instead of assuming a multi-RAT configuration, we assume that OBUs are only equipped with a single RAT, namely IEEE 802.11p. At the same time, it is also assumed that a large density of RSUs exist, with overlapping coverage areas, to avoid the service interruption during the handover process between one RSU to another. To allow the simultaneous communication between a given OBU and multiple RSUs, from now on denoted as logical links, we explore a Software Defined Vehicular Network (SDVN) responsible for the vehicular network management, and configure the MPTCP to operate on top of it. The results, obtained through experiments using real vehicular networking hardware (OBUs and RSUs) with emulated mobility, show that MPTCP can be successfully used in single-RAT environments where the radio resource is shared between the several active logical links, which are then explored by the different MPTCP subflows. To the best of our knowledge this is the first time that MPTCP is studied in single-RAT scenario, which can be considered as a paradox given the nature of such transport protocol.
@@ II SDN-based Vehicular Network
The explored Software-Defined Vehicular Network (SDVN), detailed in [7] and simplified in Figure 1, consists of OBUs, representing the mobile nodes — the vehicles — equipped with a single RAT, IEEE 802.11p with an ITS-G5 vehicular stack, and RSUs, acting as gateways to extend the connectivity of the OBUs, namely, to the Internet. The static infrastructure operates within an SDN domain, where each RSU operates as an SDN switch and an SDN controller which oversees the flow management, ensuring that downlink traffic is properly redirected to the correct logical links between the RSU(s) and the different OBUs. The management done by the SDN controller relies on Cooperative Intelligent Transport System (C-ITS) messages exchanged between RSUs and OBUs, through the vehicular network stack Vanetza [8], providing positioning information, signal strength, and other details about each mobile node to the SDN controller. A similar management must be performed by the OBU with respect to the uplink traffic, under the responsibility of the Connection Manager (CM).
@@ III MPTCP in Single-RAT Vehicular Networks
To fully explore the operation of MPTCP, the vehicular network must support the simultaneous transmission of IP traffic through different logical links between one OBU and the different RSUs in its range. One solution lies on the use of GRE tunnels in the wireless link between the OBU and the RSU, creating an overlay network [^1] It is important to highlight that this work does not aim to evaluate the performance of such overlay solution, nor the SDVN. Instead, it provides the required configuration for the proper implementation of MPTCP subflows over the several wireless logical links between the OBU and the RSUs. We are aware that other tunneling configurations could be used, such as VXLANs. [9]. In this way, the OBU’s wireless radio interface is shared by multiple GRE tunnels, associated with several RSUs, on which MPTCP subflows will be created and removed, as illustrated in Figure 2.
@@ IV Performance Evaluation
The MPTCP version used in this study is MPTCPv1222MultiPath TCP Project available at https://www.mptcp.dev/., which is the latest version and, at the time, the only one supported by the upstream Linux kernel. Furthermore, the default MPTCP scheduling algorithm (minRTT) is used, along with the Linked Increase Algorithm coupled congestion control algorithm, which couples the behavior of multiple paths [10].
@@
Although we have used a real communication vehicular network, the mobility of the OBU was emulated, meaning that the existence of a connection between the OBU and the different RSUs was done in controlled environment — turning the RSUs on and off. The OBU’s emulated path, following a straight line, is displayed in Figure 3: at $t=0$ s, the OBU is in communication range with a single RSU, RSU1; then, at $t=20$ s, the OBU becomes in range of RSU2, creating two logical communication links with RSU1 and RSU2, a situation that remains until $t=50$ s; until $t=100$ s, the OBU is in communication range of a single RSU, RSU2, and from that moment on until the end of the experimentation, the OBU is in communication range of two RSUs, RSU2 and RSU3.
@@
Figure 4(a) presents the bitrate achieved by each MPTCP subflow during the experiment for the first scenario, i.e. with no additional delay. To better understand the results achieved by the MPTCP, Figure 4(b) plots the measured RTT on each MPTCP subflow. The first observation goes to the correct use of all the available logical links between the OBU and the infrastructure: when a new RSU is in communication range, the subflow is created, and when a given RSU becomes unreachable, the respective MPTCP subflow is removed. However, the creation and removal of a MPTCP subflow does not imply a correct exploration of MPTCP: a subflow may be created but remains unused. Nevertheless, the results show that all the MPTCP subflows present considerable bitrate values.
@@
A more detailed analysis shows that the first MPTCP subflow (denoted by the MPTCP as the main subflow), which is responsible for initiating the MPTCP connection, is established at the beginning of the experiment, and data is transmitted through it, reaching an average bitrate of 8.5 Mbps [^4] All IEEE 802.11p interfaces were configured to achieve a maximum bitrate of 9 Mbps.. When the second RSU becomes in range, an additional MPTCP subflow is created, leading to a reduction of the bitrate in the first subflow by approximately half. This occurs because the capacity of the OBU’s radio interface is now shared between the two MPTCP subflows. When the connection with the first RSU is lost, we can observe an increase in the bitrate of the second MPTCP subflow justified by the fact that the radio interface is not being shared anymore between multiple logical links. The achieved bitrate is in line with the one achieved solely by the first MPTCP subflow. Finally, with the appearance of a third RSU, and consequently, a new logical link, the MPTCP explores both MPTCP subflows simultaneously.
@@ V Conclusions
This paper evaluated MultiPath TCP in a multitude of wireless network scenarios sharing a common characteristic: the existence of different radio access networks. Such configuration allows MPTCP to create a subflow over each radio access technology, allowing the simultaneous transmission of traffic through the different paths, aiming to increase the performance of the traditional TCP. This pioneering work presented a different study, answering the following questions: can we use the MPTCP in single-RAT scenarios? What is its advantage?
With a vehicular network as the use case, the results have shown that MPTCP can be effectively used when a single radio access network interface is shared to create different communication paths. According to the statistics of each wireless logical link, the MPTCP adjusts the transmission rate on each subflow, guaranteeing a proper operation with one or more active links. Such behavior provides a seamless handover process in vehicular networks, even when OBUs are equipped with a single radio access network. Future work will address multiple RATs.

### v2

# Multipath TCP with Single Radio Access Technologies: a Paradox or an Opportunity?

@@ Abstract
This paper addresses the use of Multipath Transmission Control Protocol (MPTCP) in a single Radio Access Technology (RAT) network. Different from other studies where multiple RATs are explored by the MPTCP, a situation that cannot be always guaranteed, due to lack of coverage for example, in this work we assess and evaluate the capability of MPTCP to operate over a single RAT environment. With a vehicular network as use case, we show how the IEEE 802.11p interface is shared among the multiple logical links created between the On-Board Unit (OBU) and the several Road Side Units (RSUs) in its range, supporting the different MPTCP subflows. The results, obtained through experimentation with real vehicular networking hardware, show that MPTCP allows for seamless handovers, ensuring continuous, stable and efficient communication in highly mobile environments.
@@
Vehicular Networks, Multipath Transmission Control Protocol, Single Radio Access Technologies, IEEE 802.11p.
@@ I Introduction
Multipath TCP (MPTCP) is an enhanced version of Transmission Control Protocol (TCP), designed to increase application performance and reliability by allowing concurrent use of multiple paths between the source and the destination [1]. A fundamental aspect of MPTCP is its resilience in scenarios of heterogeneous access technologies, exploiting the diversity of access technologies, dynamically distributing data across multiple paths. This ensures efficient and robust communication, allowing the protocol to adapt to the various conditions and capabilities of each network. This protocol also ensures seamless handovers between various network interfaces, such as Wi-Fi and cellular networks, a feature particularly beneficial in mobile environments, characterized by frequent changes in the network topology.
@@
This work presents a different study on the suitability and performance of MPTCP in VANETs. First, we consider the scenario where, temporarily or permanently, a single-RAT is available for the communication between OBUs and RSUs, for example, due to implementation decisions or lack of connectivity on the second RAT. Second, we consider a vehicular network architecture, enabled by Software Defined Network (SDN), where, even in a single-RAT situation, an OBU is capable of creating simultaneous logical links with different RSUs in its communication range. This characteristic allows MPTCP to operate even in single-RAT scenarios. To better understand the novelty of this study, note that operating MPTCP in this single-RAT context differs from using traditional TCP in a single-RAT scenario, as MPTCP leverages multiple distinct paths between OBUs and RSUs within its range, unlike traditional TCP, which relies on a single path.
The results, obtained through experiments using real vehicular networking hardware (OBUs and RSUs) with emulated mobility, show that MPTCP can be successfully used in single-RAT environments where the radio resources are shared between the several active logical links, which are then explored by the different MPTCP subflows. To the best of our knowledge, this is the first time that MPTCP is studied in single-RAT scenarios, which can be considered as a paradox given the nature of such transport protocol.
@@ II SDN-based Vehicular Network
The explored SDVN, detailed in [7] and simplified in Fig. 1, consists of OBUs, representing the mobile nodes — the vehicles — where a single RAT is available, namely IEEE 802.11p with an ITS-G5 vehicular stack, and RSUs, acting as gateways to extend the connectivity of the OBUs, namely, to the Internet. The static infrastructure operates within an SDN domain, where each RSU operates as an SDN switch and an SDN controller which oversees the flow management, ensuring that downlink traffic is properly redirected to the correct logical links between the RSU(s) and the different OBUs. The management done by the SDN controller relies on Cooperative Intelligent Transport System (C-ITS) messages exchanged between RSUs and OBUs, through the vehicular network stack Vanetza [8], providing positioning information, signal strength, and other details about each mobile node to the SDN controller. A similar management must be performed by the OBU with respect to the uplink traffic, under the responsibility of the Connection Manager (CM).
@@ III MPTCP in Single-RAT Vehicular Networks
To fully explore the operation of MPTCP, the vehicular network must support the simultaneous transmission of IP traffic through different logical links between one OBU and the different RSUs in its range. One solution lies on the use of GRE tunnels in the wireless link between the OBU and the RSU, creating an overlay network [^1] It is important to highlight that this work does not aim to evaluate the performance of such overlay solution, nor the SDVN. Instead, it provides the required configuration for the proper implementation of MPTCP subflows over the several wireless logical links between the OBU and the RSUs. We are aware that other tunneling configurations could be used, such as VXLANs. [9]. In this way, the OBU’s wireless radio interface is shared by multiple GRE tunnels, associated with several RSUs, on which MPTCP subflows will be created and removed, as illustrated in Fig. 2.
@@ IV Performance Evaluation
The MPTCP version used in this study is MPTCPv1222Multipath TCP Project available at https://www.mptcp.dev/., which is the latest version and, at the time, the only one supported by the upstream Linux kernel. Furthermore, the default MPTCP scheduling algorithm (minRTT) is used – where data are first transmitted in the subflow with the lowest RTT until its congestion window is full, and then it starts sending data on the subflow with the next highest RTT, and so on –, along with the Linked Increase Algorithm coupled congestion control algorithm, which couples the behavior of multiple paths [10].
@@
Although we have used a real communication vehicular network, the mobility of the OBU was emulated, meaning that the existence of a connection between the OBU and the different RSUs was done in controlled environment — turning the RSUs on and off. The OBU’s emulated path, following a straight line, is displayed in Fig. 3: at $t=0$ s, the OBU is in communication range with a single RSU, RSU1; then, at $t=20$ s, the OBU becomes in range of RSU2, creating two logical communication links with RSU1 and RSU2, a situation that remains until $t=50$ s; until $t=100$ s, the OBU is in communication range of a single RSU, RSU2, and from that moment on until the end of the experimentation, the OBU is in communication range of two RSUs, RSU2 and RSU3.
@@
Figure 4(a) presents the bitrate achieved by each MPTCP subflow during the experiment for the first scenario, i.e. with no additional delay. To better understand the results achieved by the MPTCP, Fig. 4(b) plots the measured RTT on each MPTCP subflow. The first observation goes to the correct use of all the available logical links between the OBU and the infrastructure: when a new RSU is in communication range, the subflow is created, and when a given RSU becomes unreachable, the respective MPTCP subflow is removed. However, the creation and removal of a MPTCP subflow does not imply a correct exploration of MPTCP: a subflow may be created but remains unused. Nevertheless, the results show that all the MPTCP subflows present considerable bitrate values.
@@
A more detailed analysis shows that the first MPTCP subflow (denoted by the MPTCP as the main subflow), which is responsible for initiating the MPTCP connection, is established at the beginning of the experiment, and data is transmitted through it, reaching an average bitrate of 8.5 Mbps [^4] IEEE 802.11p radios were configured for a maximum bitrate of 9 Mbps.. When the second RSU becomes in range, an additional MPTCP subflow is created, leading to a reduction of the bitrate in the first subflow by approximately half. This occurs because the capacity of the OBU’s radio interface is now shared between the two MPTCP subflows. When the connection with the first RSU is lost, we can observe an increase in the bitrate of the second MPTCP subflow justified by the fact that the radio interface is not being shared anymore between multiple logical links. The achieved bitrate is in line with the one achieved solely by the first MPTCP subflow. Finally, with the appearance of a third RSU, and consequently, a new logical link, the MPTCP explores both MPTCP subflows simultaneously.
@@ V Conclusions
In mobile networks, the connectivity between the mobile nodes and the infrastructure is pivotal: without it, the mobile nodes may become isolated. To reduce such risk, mobile nodes are usually equipped with multiple RATs, complementing each other. Such configuration is explored by the MPTCP, where each MPTCP subflow is established over each RAT, allowing the simultaneous transmission of traffic through the different paths, with the purpose of increasing the performance of the traditional TCP. However, assuming that such multi-connectivity is always available is far from the reality. Thus, this pioneering work presents a different perspective on the use of MPTCP: can we use it in single-RAT environments?
With a vehicular network as use case, the results have shown that MPTCP can be effectively used when a single radio access technology is shared to create simultaneous communication paths between the mobile node and the infrastructure. Based on the quality of each wireless logical link, the MPTCP adjusts the transmission rate on each subflow, guaranteeing a proper operation with one or more active links. Such behavior provides a seamless handover process in vehicular networks, even when OBUs are equipped with a single radio access technology. Future work will extend the single-RAT configuration to include additional RATs to support more MPTCP subflows.